Activity 27 continued Lead-210 also has a radioactive decay rate. Its half-life is 22 years. What is k, its exact decay rate? If we assume that we begin with an initial amount of of lead, write a function that approximately models the lead-210 at time, t, in years. significant figures for the function m(t) Fill in the t-values and answer the questions with function equations. Include units. Be exact. It might be easier without your calculator. Exactly how much will remain after 44 years? m(. Exactly how much would there have been 22 years ago? Exactly how much will remain after 110 years? When will there be 500g of lead? Include units. When will there be 125g of lead? Include units. When was there 4000g of lead? Include units. Graph the function 4000 2000 20 40 60 80 100 120 20 40 2000 ndo on environmental conditions
Answer #1

Half-life t1/2 = 22 years

t1/2 = 0.6931/k

k = 0.6931/22

k= 0.0315

The function of decay

N = Noe-kt

After 44 years

Here No=1000 gram and k = 0.032

N = 1000e-0.032x44

N = 244.63 grams

22 years ago

Here N=1000 gram and k = 0.032

1000= Noe-0.032x22

N0 = 2021.84 grams

After 110 years

Here No=1000 gram and k = 0.032

N = 1000e-0.032x110

N = 29.6 grams

To find t for 500 gram

500 = 1000 e-0.032xt

0.5 = e-0.032t

ln 0.5 = -0.032t

t = 21.66 years

to find t for 125 g

125 = 1000 e-0.032xt

0.125 = e-0.032t

ln 0.125 = -0.032t

t = 64.98 years

.. 4000 2000 . -40 -20 20 40 60 80 100 120 ..-2000....